# Generalizable data-driven turbulence closure modeling on unstructured grids with differentiable physics

### v2

@@ 3 Datasets and training
We optimize the ML model by minimizing the $L^{2}$ norm error of the resulting a-posteriori velocity field with respect to the training dataset from the ground truth velocity field. During training, a predicted solution trajectory $\hat{\mathbf{u}}$ is obtained by simulating the LES equations using the GNN-based SGS closure model. The training dataset is generated using only the first 256 snapshots of the LES-80k simulation. We set multiple trajectories for training dataset, where each of trajectory consists of 32 consecutive snapshots with different initial conditions separated by 16 timesteps. This is based on our own empirical analysis and suggestions from previous work to avoid gradient explosion. Ultimately, this leads to a total of 15 trajectories for training. The velocity field in each trajectory $\{\mathbf{u}_{k}\}_{k=0}^{n}$ is projected onto the coarse mesh used in the ML-LES simulation to generate the target velocity field $\{\overline{\mathbf{u}}_{k}\}_{k=0}^{n}$. The projection is performed using the Galerkin projection method as follows. In a FEM-based CFD solver, the flow field data is represented as a function of the $x$, $y$ and $z$ coordinates within a specified function space. In this study, the two-dimensional ground truth velocity field, $\mathbf{u}$, defined on the dense grid, is represented within each cell as a function of $x$ and $y$ and belongs to a second-order continuous Galerkin function space. Likewise, the ground truth velocity data projected onto the coarse grid, $\overline{\mathbf{u}}$, is defined on each cell of the coarse grid and belongs to a second-order continuous Galerkin function space based on the coarse mesh. The weak form equation of the projection from $\mathbf{u}$ to $\overline{\mathbf{u}}$ is expressed as follows:
[…]
@@
In FEniCS, a second-order Lagrange function is chosen as the basis function within the second-order continuous Galerkin function space, where each component of $\mathbf{c_{e}}$ represents the coefficient of a basis function for $\overline{\mathbf{u}}$. Since the linear system in Eq. 28 is defined locally on each element, the global linear system is assembled to ensure continuity between elements as follows.
@@
This assembly process combines the contributions of each element, resulting in a system that enforces the continuity of the solution across shared boundaries between adjacent elements. The pressure field is also projected for the initial condition. The initial condition is integrated over 32 steps using the GNN-based turbulence model. The overall loss is computed as
@@
where both the instantaneous and the time-averaged accuracy are considered, and the integral is computed using the FEM discretization. This formulation of the loss function enables the model to learn sub-grid closures that not only match the instantaneous flow field values but also maintain consistency with the overall trend of actual values. Since the integral is computed in FEniCS, an adjoint can be obtained using the same method described previously. The loss is backpropagated in an end-to-end fashion, through all PDE evaluations and timesteps, to update the parameters of the GNN. The model is trained for 50 epochs using the Adam optimizer [75] with a learning rate of $10^{-4}$. Training took about 6 hours on 2 NVIDIA A100 GPUs and 8 AMD EPYC 7543P CPU-cores. Much of the inefficiency in the model can be attributed to data transfer between the CPU and GPU, where the CPU is used to solve the PDE and the GPU is used to evaluate the GNN. Training history is shown in the Fig. 23 in Appendix. A.
4 Results and discussion
@@
[…]
4.1 Training result
We first examine the results from the BFS case. Fig. 4 shows the normalized $L^{2}$ error over time for the four models tested (with the gray dashed line representing the extent of training data observed). We observe that NM-5k shows the steepest growth and highest overall error, indicating the need for a closure model. LES-5k displays a smaller initial slope and fluctuates around a significantly lower error than the NM-5k model, while the LES-20k model results in the smallest initial slope and similar stationary error as LES-5k. We see that the learned model ML-5k generally results in the lowest overall $L^{2}$ error, and given that this metric is reflective of the training cost function, we determine that the training process is functioning as intended to minimize the $L^{2}$ error. Furthermore, the ML-5k model is stable for longer time integration than time steps used during training, $t=0.0256s$. However, the transient $L^{2}$ error is just one metric we should use to evaluate model performance. To further examine the transient behavior of the model, pointwise accuracy of each model near the wall is compared. Fig. 5 illustrates pointwise accuracy of the models from a probe at $(x,y)=(h,h)$ using the normalized x-direction velocity. It is observed that the ML-5k model is more accurate in capturing periodic fluctuations of velocity than other models and remains stable for longer time horizon.
Figure 4: Normalized $L_{2}$ error on the BFS case. The gray dashed line represents the extent of training data collected.
@@ 4.1 Training result
The mean flow is a particularly useful measure that decouples the transient effects of the model from flow features of interest. Fig. 6 depicts the normalized x-velocity of the flow and the resulting error field with respect to the LES-80k model. We also report the integrated total error for each model. We can observe the recirculation region characterized by the negative (blue) x-velocity that is present in each model. the NM-5k model contains discrepancies throughout the portion of the domain shown, including prior to the step, in front of the step, and at the boundary layer. The LES-5k model reduces the error by a factor of over 5, but still contains deviations, notably at the boundary layer, while the LES-20k model achieves smaller error than the LES-5k model by reducing the error in this region. The ML-5k model achieves the smallest errors among models by reducing the error within the region in front of the step.
The outputs of our sub-grid closure model include the Smagorinsky coefficient, $C_{s}$, and the coefficients of four independent basis matrices of the non-linear sub-grid stress tensor, $k$, $\alpha$, $\beta$, and $\gamma$. The final non-linear stress tensor can be constructed by summing these basis matrices in the following manner:
$\bm{\tau}^{NL}$ $=\begin{bmatrix}\tau^{NL}_{xx}&\tau^{NL}_{xy}\\ \tau^{NL}_{yx}&\tau^{NL}_{yy}\end{bmatrix}=k\begin{bmatrix}1&0\\ 0&1\end{bmatrix}+\alpha\begin{bmatrix}1&0\\ 0&-1\end{bmatrix}+\beta\begin{bmatrix}0&1\\ 1&0\end{bmatrix}+\gamma\begin{bmatrix}0&1\\ -1&0\end{bmatrix}=\begin{bmatrix}k+\alpha&\beta+\gamma\\ \beta-\gamma&k-\alpha\end{bmatrix}$ (32)
@@
Consequently, each element of the non-linear sub-grid stress tensor, $\tau^{NL}_{ij}$, can be visualized as the sum or difference of the coefficients of the basis matrices. Fig. 7 illustrates the visualization of the normalized x-direction velocity, the Smagorinsky coefficient, and each component of the non-linear sub-grid stress tensor predicted by the sub-grid closure model at $t=0.02s$. In LES simulation with the conventional Smagorinsky model, $C_{s}$, is predetermined before the simulation to model the viscous effect of sub-grid scale eddies, and every mesh point in the computational domain has the same $C_{s}$ value, which remains fixed throughout the simulation. This often leads to excessive dissipation in the viscous sub-layer region near walls where wall viscosity is already strong enough to damp the eddies. Furthermore, flow state predictions near vortices with high local shear can be inaccurate due to this over-dissipation. In contrast, the ML-5k adaptively adjusts $C_{s}$ during the simulation to tune the sub-grid dissipation. Specifically, the ML-5k model predicts smaller $C_{s}$ values in the near-wall regions compared to the areas farther from the walls. In addition, the model predicts lower $C_{s}$ value near the vortices where local shear is high due to rapid velocity changes around the vortex cores. This adaptive adjustment of $C_{s}$ reduces the dissipation caused by the sub-grid eddies, thereby decreasing errors within the near-wall region and region upstream of the step as shown in Fig. 6. We also observe that the non-linear stress components exhibit large absolute values near the separation region and vortices, where the Smagorinsky coefficient is small and, consequently, the stress predicted by the linear eddy viscosity model is also small. This indicates that our model can effectively identify regions with high values or sudden changes in strain rate, where the Boussinesq approximation [76] of linear eddy viscosity is challenged. In addition, the shear stress terms, $\tau^{NL}_{xy}$ and $\tau^{NL}_{yx}$, are predicted to exhibit much larger magnitude in the negative direction compared to the normal stress terms, $\tau^{NL}_{xx}$ and $\tau^{NL}_{yy}$, near vortices. This demonstrates that our model effectively responds to the non-linear and complex turbulent interactions around vortices, including the non-linear momentum exchange between the $x$- and $y$-direction velocity components and backscatter. It is notable that such complex sub-grid learning is extracted despite our model being exposed to a training dataset solely comprised of sampled velocity time-series across the domain.
Figure 7: Visualization of the normalized X-direction velocity, Smagorinsky coefficient, and each component of the non-linear sub-grid stress tensor at $t=0.02sec$
4.2 Generalization to new geometries
We apply our GNN-based subgrid closure model, trained on the BFS case, to different geometries to evaluate its generalization ability. These new geometries include ramp and wall-mounted cube (WMC) cases with the same height of $h=0.0127m$ as the BFS case and are tested under the same Reynolds number, $Re=26,501$. They present a more challenging extrapolation task for the trained model than simply changing the step size in the BFS case, as these configurations have distinct geometric differences near the separation regions and exhibit different separation physics. Visualizations of the discretizations of these geometries are provided in Fig.24 in Appendix. B
Fig. 8 shows the normalized $L^{2}$ error over time for the ramp case, where we compare the performance of the trained model with three other baseline predictions. We observe that NM-5k exhibits the highest overall error (due to the lack of any closure), while LES-5k has lower error than NM-5k, highlighting the need for a closure model. Furthermore, ML-5k shows a lower error than LES-5k, whereas LES-20k achieves the lowest error. This suggests that even when tested on the ramp case, which has a smoother separation due to a smaller flow turning angle, our model is improved compared to LES-5k model. Furthermore, the proposed model exhibits stability beyond the time steps used during training, suggesting its robustness in extended time-step integration. Fig. 9 presents the pointwise accuracy of the models based on normalized x-direction velocity data obtained from a probe located at $(x,y)=(2h,h)$. The trends observed in the pointwise analysis are consistent with the overall error profiles, demonstrating that the ML model not only improves upon the performance of LES-5k but also remains stable for longer prediction horizons than those encountered during training.
Figure 8: Normalized $L_{2}$ error on the ramp case. This geometry was unseen during training.
Figure 9: X-direction Velocity measurement taken from a probe in front of the ramp at $(x,y)=(2h,h)$
Fig. 10 presents a contour plot of the normalized mean x-direction velocity around the ramp, along with the error distribution for each model when compared to the LES-80k reference solution. Additionally, the integrated error values for each model are reported. Across all models, the recirculation zones immediately downstream of the ramp are well captured. However, the NM-5K and LES-5k models exhibit the largest errors near these recirculation zones and along the wall boundary. In contrast, the LES-20k model shows relatively smaller errors in these areas, achieving the lowest error among the models. The ML-5k model effectively reduces errors in the high shear flow and near-wall turbulence regions by adaptively modulating the Smagorinsky coefficient. Fig. 11 shows a visualization of the Smagorinsky coefficient and the normalized x-direction velocity across the domain at $t=0.02s$. It is observed that our model changes $C_{s}$ dynamically depending on localized flow quantities, thereby reducing errors near the wall boundary and vortices.
Figure 11: (Top row) Smagorinsky coefficient and (Bottom row) Normalized X-direction velocity predicted by trained model across domain near the ramp and their zoomed-in views at t=0.02
Fig. 12 compares the predictability of the trained model for the WMC case with three other models using the $L^{2}$ error over time. In the WMC case, simulations without a turbulence model become unstable and diverge after several time steps, indicating the necessity of incorporating a closure model. While the LES-20k achieves the lowest error among the models, using the ML-5k model does not result in significant performance improvement over the LES-5k model unlike the ramp case. This can be attributed to the difference in separation mechanisms between two cases. In contrast to the ramp case where flow originates from inflow and simply separate near the slope, the flow upstream of the cube encounters a stagnation point with high pressure and recirculation before separation—a fundamentally different separation mechanism from what the trained model is exposed to during training. Nevertheless, the trained model recovers the performance of the LES-5k model and remains stable in extended time-step integration above time span used during training. Fig. 13 compares the pointwise accuracy of models from a probe at $(x,y)=(6h,3.5h)$ using the x-direction velocity normalized by the freestream velocity. At this location, the LES-20k model can capture peaks of the periodic fluctuation of the true data, while the LES-5k model predicts lower peaks than the actual values. Although using the trained model does not provide a marked improvement in peak prediction, ML-5k remains stable for longer predictions horizons than those used during training.
Figure 12: Normalized $L_{2}$ error on the WMC case. This geometry was unseen during training.
Figure 13: X-direction Velocity measurement taken from a probe in front of the cube at $(x,y)=(6h,3.5h)$
Fig. 14 illustrates the normalized mean x-direction velocity field and the corresponding error distribution around the cube, along with the integrated error values in comparison with the LES-80k model. Across all models, the recirculation zones upstream and downstream of the cube are effectively captured. However, the LES-5k model exhibits large errors near the vortex trajectories and wall boundaries. In contrast, the LES-20k model shows significantly smaller errors in these regions. The ML-5k model achieves a lower integrated error compared to the LES-5k model by reducing errors near these areas. Fig. 15 illustrates adaptive adjustment of the Smagorinsky coefficient and the normalized x-direction velocity across domain. This figures shows that the ML-5k model decreases $C_{s}$ value near the vortex-formulating region and the wall boundary layer.
@@ 4.2 Generalization to new geometries
Figure 14: Plots of the mean flow from the WMC case for various models. (Left column) The normalized mean x-velocity and (Right column) The error with respect to the truth in space, as well as the integrated total.
@@
Figure 15: (Top row) Smagorinsky coefficient and (Bottom row) Normalized X-direction velocity predicted by trained model across domain near the cube and their zoomed-in views at t=0.02

### v3

@@ III.1.2 Training procedure and loss formulation
We optimize the ML model by minimizing the $L^{2}$ norm error of the resulting a-posteriori velocity field with respect to the training dataset from the ground truth velocity field. During training, a predicted solution trajectory $\hat{\mathbf{u}}$ is obtained by simulating the LES equations using the GNN-based SGS closure model. The training dataset is generated using only the first 256 snapshots of the LES-80k simulation. We set multiple trajectories for training dataset, where each of the trajectories consists of 32 consecutive snapshots with different initial conditions separated by 16 timesteps. This is based on our own empirical analysis and suggestions from previous work to avoid gradient explosion [49]. Ultimately, this leads to a total of 15 trajectories for training. The velocity field in each trajectory $\{\mathbf{u}_{k}\}_{k=0}^{n}$ is projected onto the coarse mesh used in the ML-LES simulation to generate the target velocity field $\{\overline{\mathbf{u}}_{k}\}_{k=0}^{n}$. The projection is performed using the Galerkin projection method as follows. In a FEM-based CFD solver, the flow field data is represented as a function of the $x$, $y$ and $z$ coordinates within a specified function space. In this section, the two-dimensional ground truth velocity field, $\mathbf{u}$, defined on the dense grid, is represented within each cell as a function of $x$ and $y$ and belongs to a second-order continuous Galerkin function space. Likewise, the ground truth velocity data projected onto the coarse grid, $\overline{\mathbf{u}}$, is defined on each cell of the coarse grid and belongs to a second-order continuous Galerkin function space based on the coarse mesh. The weak form equation of the projection from $\mathbf{u}$ to $\overline{\mathbf{u}}$ is expressed as follows:
$\int_{\Omega_{e}}\overline{\mathbf{u}}\cdot v\;dx=\int_{\Omega_{e}}\mathbf{u}\cdot v\;dx,\quad\forall v\in{V_{c}},$ (24)
@@
In FEniCS/FEniCSx, a second-order Lagrange function is chosen as the basis function within the second-order continuous Galerkin function space, where each component of $\mathbf{c_{e}}$ represents the coefficient of a basis function for $\overline{\mathbf{u}}$. Since the linear system in Eq. 26 is defined locally on each element, the global linear system is assembled to ensure continuity between elements as follows.
@@
This assembly process combines the contributions of each element, resulting in a system that enforces the continuity of the solution across shared boundaries between adjacent elements. The pressure field is also projected for the initial condition. The initial condition is then integrated over 32 steps using the GNN-based turbulence model. The overall loss is computed as
@@
where both the instantaneous and the time-averaged accuracy are considered, and the integral is computed using the FEM discretization. This formulation of the loss function enables the model to learn subgrid closures that not only match the instantaneous flow field values but also maintain consistency with the overall trend of actual values. Since the integral is computed in FEniCS an adjoint can be obtained using the same method described in Sec. II.3. The loss is backpropagated in an end-to-end fashion, through all PDE evaluations and timesteps, to update the parameters of the GNN. The model is trained for 50 epochs using the Adam optimizer [85] with a learning rate of $10^{-4}$. Training took about 6 hours on 2 NVIDIA A100 GPUs and 8 AMD EPYC 7543P CPU-cores. Much of the inefficiency in the model can be attributed to data transfer between the CPU and GPU, where the CPU is used to solve the PDE and the GPU is used to evaluate the GNN. Training history is shown in the Fig. 43 in Appendix. A.
III.1.3 Training results for 2D flows
@@
where $U_{\infty}^{2}$ is the inlet velocity, and $\Omega$ is the computational domain. Additionally, we provide the velocity sampled from a probe in front of the step at the location $(h,h)$. This location is based on the region where the error between the LES-5k and the LES-80k models is the largest. For these transient plots, we report values from the first 1024 timesteps, extending well beyond the training dataset, which consists of the first 256 timesteps. This allows us to evaluate the accuracy and stability of pointwise predictions over a longer prediction horizon. We also examine the time-averaged statistics of the flow, as these are often the most relevant metrics of interest. We present the time-averaged x-velocity field $U_{x}$ along with the corresponding error field to assess each model’s ability to accurately capture the recirculation zone and identify areas of discrepancy within the flow. This series of a-posteriori metrics provides a comprehensive evaluation of model performance, ensuring that the predicted flow aligns closely with the ground truth.
Figure 4 shows the normalized $L^{2}$ error over time for the four models tested (with the gray dashed line representing the extent of training data used). We observe that NM-5k shows the steepest growth and highest overall error, indicating the need for a closure model. LES-5k shows a smaller initial rate of error increase and fluctuates around a significantly lower error than the NM-5k model, while the LES-20k model results in the smallest initial rate of error increase and smaller error than LES-5k. We see that the learned model ML-5k generally results in the lowest overall $L^{2}$ error, and given that this metric is reflective of the training cost function, we determine that the training process is functioning as intended to minimize the $L^{2}$ error. Furthermore, the ML-5k model is stable for more time steps than those used during training ($t=0.0256s$). However, the transient $L^{2}$ error is just one metric we should use to evaluate the trained model’s performance. To further examine the transient behavior of the model, pointwise accuracy of each model near the wall is compared. Fig. 5 illustrates pointwise accuracy of the models from a probe at $(x,y)=(h,h)$ using the normalized $x$-direction velocity. It is observed that the ML-5k model is more accurate in capturing periodic fluctuations of velocity than other models and remains stable for a longer time horizon.
Figure 4: Normalized $L_{2}$ error on the two-dimensional BFS case. The gray dashed line represents the extent of training data collected.
@@
The mean flow is a particularly useful measure that decouples the transient effects of the model from flow features of interest. Fig. 6 depicts the normalized x-velocity of the flow and the resulting error field with respect to the LES-80k model. We also report the integrated total error for each model. We can observe the recirculation region characterized by the negative (blue) x-velocity that is present in each model. the NM-5k model contains discrepancies throughout the portion of the domain shown, including prior to the step, in front of the step, and at the boundary layer. The LES-5k model reduces the error by a factor of over 5, but still contains deviations, notably at the boundary layer, while the LES-20k model achieves smaller error than the LES-5k model by reducing the error in this region. The ML-5k model achieves the smallest errors among models by reducing the error within the region downstream of the step.
@@
Figure 7 illustrates the visualization of the normalized $x$-direction velocity and the Smagorinsky coefficient predicted by the subgrid closure model at $t=0.02s$. In LES simulation with the conventional Smagorinsky model, $C_{s}$ is predetermined before the simulation to model the viscous effect of subgrid-scale eddies, and every mesh point in the computational domain has the same $C_{s}$ value, which remains fixed throughout the simulation. This often leads to excessive dissipation in the viscous sub-layer region near walls where wall viscosity is already strong enough to damp the eddies. Furthermore, flow state predictions near vortices with high local shear can be inaccurate due to this over-dissipation. In contrast, the ML-5k adaptively adjusts $C_{s}$ during the simulation to tune the subgrid dissipation. This is possible because our ML subgrid closure model can learn the optimal amount of local subgrid dissipation to minimize the ground truth error. Specifically, the ML-5k model predicts smaller $C_{s}$ values in the near-wall regions compared to the areas farther from the walls. In addition, the model predicts lower $C_{s}$ value near the vortices where local shear is high due to rapid velocity changes around the vortex cores. This adaptive adjustment of $C_{s}$ reduces the dissipation caused by the subgrid eddies, thereby decreasing errors within the near-wall region and the region upstream of the step as shown in Fig. 6. It is notable that such complex subgrid learning is extracted despite our model being exposed to a training dataset solely comprised of sampled velocity time-series across the domain.
@@
Figure 6: Plots of the mean flow from two-dimensional BFS case for various models. (Left column) The normalized mean x-velocity and (Right column) The error with respect to the truth in space, as well as the integrated total.
@@
Figure 7: (Top row) Smagorinsky coefficient and (Bottom row) Normalized $x$-direction velocity predicted by trained model across domain near the two-dimensional BFS and their zoomed-in views at t=0.02
III.1.4 Generalization to new geometries
We apply our GNN-based subgrid closure model, trained on the two-dimensional BFS case, to different geometries to evaluate its generalization ability. These new geometries include two-dimensional ramp and wall-mounted cube (WMC) cases with the same height of $h=0.0127m$ as the BFS case and are tested under the same Reynolds number, $Re=26,501$. They present a more challenging extrapolation task for the trained model as these configurations have distinct geometric differences near the separation regions and exhibit different separation physics. Visualizations of the discretizations of these geometries are provided in Fig.45 in Appendix. B
Fig. 8 shows the normalized $L^{2}$ error over time for the two-dimensional ramp case, where we compare the performance of the trained model with three other baseline predictions. We observe that NM-5k exhibits the highest overall error (due to the lack of any closure), while LES-5k has lower error than NM-5k, highlighting the need for a closure model. Furthermore, ML-5k shows a lower error than LES-5k, whereas LES-20k achieves the lowest error. This suggests that even when tested on the ramp case, which has a smoother separation due to a smaller flow turning angle, our model is improved compared to LES-5k model. Furthermore, the proposed model exhibits stability beyond the time steps used during training, suggesting its robustness in extended time-step integration. Fig. 9 presents the pointwise accuracy of the models based on normalized $x$-direction velocity data obtained from a probe located at $(x,y)=(2h,h)$. The trends observed in the pointwise analysis are consistent with the overall error profiles, demonstrating that the ML model not only improves upon the performance of LES-5k but also remains stable for longer prediction horizons than those encountered during training.
Figure 8: Normalized $L_{2}$ error on the two-dimensional ramp case. This geometry was unseen during training.
Figure 9: $x$-direction Velocity measurement taken from a probe in front of the two-dimensional ramp at $(x,y)=(2h,h)$
Figure 10 presents a contour plot of the normalized mean $x$-direction velocity around the two-dimensional ramp, along with the error distribution for each model when compared to the LES-80k reference solution. Additionally, the integrated error values for each model are reported. Across all models, the recirculation zones immediately downstream of the ramp are well captured. However, the NM-5K and LES-5k models exhibit the largest errors near these recirculation zones and along the wall boundary. In contrast, the LES-20k model shows relatively smaller errors in these areas, achieving the lowest error among the models. The ML-5k model effectively reduces errors in the high shear flow and near-wall turbulence regions by adaptively modulating the Smagorinsky coefficient. Fig. 11 shows a visualization of the Smagorinsky coefficient and the normalized $x$-direction velocity across the domain at $t=0.02s$. It is observed that our model changes $C_{s}$ dynamically depending on localized flow quantities, thereby reducing errors near the wall boundary and vortices.
Figure 10: Plots of the mean flow from the two-dimensional ramp case for various models. (Left column) The normalized mean x-velocity and (Right column) The error with respect to the truth in space, as well as the integrated total.
Figure 11: (Top row) Smagorinsky coefficient and (Bottom row) Normalized $x$-direction velocity predicted by trained model across domain near the two-dimensional ramp and their zoomed-in views at t=0.02
Figure 12 compares the predictability of the trained model for the two-dimensional WMC case with three other models using the $L^{2}$ error over time. In the WMC case, a simulation without a turbulence model becomes unstable and diverges after several time steps, indicating the necessity of incorporating a closure model. While the LES-20k achieves the lowest error among the models, using the ML-5k model does not result in significant performance improvement over the LES-5k model unlike the ramp case. This can be attributed to the difference in separation mechanisms between the two cases. In contrast to the ramp case where flow originates from inflow and simply separates near the slope, the flow upstream of the cube encounters a stagnation point with high pressure and recirculation before separation—a fundamentally different separation mechanism from what the trained model is exposed to during training. Nevertheless, the trained model recovers the performance of the LES-5k model and remains stable in extended time-step integration above time span used during training. Figure 13 compares the pointwise accuracy of models from a probe at $(x,y)=(6h,3.5h)$ using the $x$-direction velocity normalized by the freestream velocity. At this location, the LES-20k model can capture peaks of the periodic fluctuation of the true data, while the LES-5k model predicts lower peaks than the actual values. Although using the trained model does not provide a marked improvement in peak prediction, ML-5k remains stable for longer prediction horizons than those used during training.
Figure 12: Normalized $L_{2}$ error on the two-dimensional WMC case. This geometry was unseen during training.
Figure 13: $x$-direction Velocity measurement taken from a probe in front of the two-dimensional cube at $(x,y)=(6h,3.5h)$
Figure 14 illustrates the normalized mean $x$-direction velocity field and the corresponding error distribution around the cube, along with the integrated error values in comparison with the LES-80k model. Across all models, the recirculation zones upstream and downstream of the cube are effectively captured. However, the LES-5k model exhibits the largest errors near the vortex trajectories and wall boundaries. In contrast, the LES-20k model shows significantly smaller errors in these regions. The ML-5k model achieves a lower integrated error compared to the LES-5k model by reducing errors near these areas. Figure 15 illustrates adaptive adjustment of the Smagorinsky coefficient and the normalized $x$-direction velocity across domain. This figures shows that the ML-5k model decreases $C_{s}$ value near the vortex-forming region and the wall boundary layer.
Figure 14: Plots of the mean flow from the two-dimensional WMC case for various models. (Left column) The normalized mean x-velocity and (Right column) The error with respect to the truth in space, as well as the integrated total
Figure 15: (Top row) Smagorinsky coefficient and (Bottom row) Normalized $x$-direction velocity predicted by trained model across domain near the two-dimensional cube and their zoomed-in views at t=0.02